OpenStudy (anonymous):

Ahmad is 5 years older than his wife who is five times as old as her daughter.if the daughter wa 4 years old 3 years back.what is the age of Ahmad?

OpenStudy (anonymous):

3 years back or Current?

OpenStudy (anonymous):

age of Ahmad is 40

OpenStudy (anonymous):

d=4+3 w=5d a=w+5 Therefore: Ahmad's age = 5(4+3)+5

OpenStudy (anonymous):

back

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

NP!! :-) And thank you! :-)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

The population of a town increases at the rate of 5 % every year.The present population is 8000.In how many years it will become 9200? please sove this one